what is the equation of a line that passes through (-4,8)and is parallel to 6x-3y=9
Answer by ewatrrr(24785)


Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept
Line parallel to 6x-3y = 9 OR y = 2x - 3 (blue line) slope m = 2
parallel line have identical slopes. thus new line:
y = 2x + b using ordered pair pt(-4,8) to solve for b
8 = 2*(-4) + b
16 = b
y = 2x + 16 Green line